40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 8 20:34:29 2012, Jenne, Update, ASS, Trouble measuring sensing matrix 
    Reply  Wed Aug 8 20:50:39 2012, Koji, Update, ASS, Trouble measuring sensing matrix 
       Reply  Thu Aug 9 00:23:11 2012, Jenne, Update, ASS, Trouble measuring sensing matrix 
          Reply  Thu Aug 9 01:26:03 2012, Koji, Update, ASS, Trouble measuring sensing matrix 
Message ID: 7129     Entry time: Thu Aug 9 00:23:11 2012     In reply to: 7124     Reply to this: 7131
Author: Jenne 
Type: Update 
Category: ASS 
Subject: Trouble measuring sensing matrix 

Quote:

From the log, I couldn't understand what has been done.

The procedure we should perform is

  1. Dither total 4 dofs of the two mirrors with different frequencies. Some fluctuation of TRY is even visible in dataviewer.
  2. The cavity is aligned at the beginning. These 4 peaks in TRY in DTT is small or invisible. Some 2nd harmonics are visible.
  3. Misalign one of the dofs. Some peaks get bigger.
  4. Correspoding lockin output becomes bigger.

Then you can start measuring the sensing matrix. At which part did the attempt fail?

Quote:
 

 Cavity started out aligned pretty well, but not 100%.  Transmission was ~0.8 . Perhaps this was part of the problem.

I realize now that you mention it, it was totally amateur hour of me to only look at the lockin outputs on StripTool (plus POY and TRY on Dataviewer), and not look at TRY on DTT...or any spectra at all.  Not so intelligent.  I could see some fluctuation of TRY on Dataviewer that corresponded to me turning on the oscillators, as well as the spot wiggling on the camera view of ETMYT a teeny bit.

When applying a 10% misalignment to ETMY Pit (by adding 0.1 to the Pit components of the output matrix, as is done in the MC spot position calibration), I could see that there was a small jump in the StripTool trace, but it was much smaller than the ambient fluctuations of the output. 

I just looked back and realized that I must have forgotten to add my screenshot, but it's saved on a desktop on Rossa.  It would be better if I had attached the data, but from that you see that the average of the lockin output signal didn't change very much in the last several minutes of the measurement, but the fluctuations (no misalignment offsets) are pretty big, maybe ~10% or 15% the size of the signal.  Then when I added the misalignment to one mirror (ETMY PIT), there is a very small jump in the lockin signal, but it is much, much smaller than the size of the ambient fluctuations.  Perhaps a long average would result in a "real" value, but by looking by eye, I can't see a discernible difference in the average value of the lockin outputs.

My plan is to do as you say, dithering all 4 optics, and misaligning a single optic's single DoF (Pit or Yaw), and seeing how that misalignment affected each of the sensors (the lockin outputs).  Then put that DoF back to nominal, and misalign a different DoF, rinse and repeat.

Okay, so this is a little stream-of-consciousness-y, and you're going to think I'm really dumb, but I just realized that I haven't set the phase of the lockin demodulators yet.  So I think I need to dither the optics, and go through each of the sensors, and adjust the phase until the peak in TRY in DTT is maximized for the I phase, and minimized for the Q phase (since we use the I-output).  Bah.  Bad Jenne.

ELOG V3.1.3-